factor the quadratic expression completely, and find the roots of the expression.
49x^2 - 14x - 3

Answer by vidhyak(98)

Solution
ax^2-bx-c
49x^2 - 14x - 3
Factorize the variable b (in our case -14) into 2 numbers (say b1,b2) so that b1* b2 = ac(49*-3) and b1+b2 = b
We can split the number -14 as -21, +7 so that (-21*7 = -147 which is equal = ac (49*-3))
and -21+7 = -14 which is b

49x^2 - 14x - 3
49x^2 - 21x + 7x - 3
7x(7x-3)+7x-3
7x(7x-3)+1(7x-3)
(7x+1)(7x-3)
Solving we get x= -1/7 and 3/7
or -.1428 and 0.4285